OpenStudy (anonymous):

A drama club is planning a bus trip to New York City to see a Broadway play. The table represents the cost per person for the bus rental compared to the number of people going on the trip. What function models the data, and how much per person will it cost if 12 students go on the trip? Number of Students(n) - Cost per Student(c) 3 - 24$ 6 - 12$ 9 - 8$ 16 - $4.5 A. n/c = 72, $12 B. nc = 9, $10 C. nc = 72, $6 D. n/c = 9, $12

OpenStudy (anonymous):

You could try to solve this by seeing it as a geometric progression, where the price per person goes down as you get more people going on the trip. Since it was multiple choice, I just inspected the answer choices to see if any seemed possible.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I noticed that the number of students multiplied by the cost... in other words, n*c... for each of the examples given in the question always gave n*c = 72. 24 * 3 = 72 6 * 12 = 72 9 * 8 = 72 16 * 4.5 = 72

OpenStudy (anonymous):

That makes it appear the answer is (c), since it's the only one where nc = 72. And to check, if you have 12 students, then n = 12, and if nc = 72, then the cost should be 72/12 = 6 dollars.

OpenStudy (anonymous):

so the answer is c..

OpenStudy (anonymous):

or are you saying its not c

OpenStudy (anonymous):

no, I think it is c... I was just showing that both parts of the answer seem to make sense. nc = 72, for n = 12 students and c = $6 per student

OpenStudy (anonymous):

o ok I was a little confused at first lol

OpenStudy (anonymous):

Yes, the way this is worded is a little weird. But it is "c". You might eventually figure out that c = 72 / n, so for any number of students "n", you find the cost by dividing 72 by that number "n". But that's not an obvious cost model... sort of a "group rate" that changes continually with every additional student.

OpenStudy (anonymous):

ok i understand now...thank you so much for your help and time

OpenStudy (anonymous):

glad to help :) Good luck!

OpenStudy (anonymous):

thank you